OpenStudy (anonymous):

Recall from Example 1 that whenever Suzan sees a bag of marbles, she grabs a handful at random. She has seen a bag containing three red marbles, three green ones, four white ones, and two purple ones. She grabs seven of them. Find the probability of the following event, expressing it as a fraction in lowest terms. HINT [See Example 1.] She has at least one green one.

OpenStudy (kropot72):

The probability of there being no green marbles in the sample of 7 is given by: \[P(0\ green)=\frac{9C7}{12C7}=\frac{9!\times7!\times5!}{2\times7!\times12!}=you\ can\ calculate\] The probability of at least one green is: 1 - P(0 green) = you can calculate.